Box 11.3 Supplement Worked example Imagine it is necessary to produce a tablet weighing 50 mg which contains 75 μg of a potent steroid, and that the product specification requires 99.5% of tablets to contain between 49 ug and 51 ug of the steroid. If the mean particle density of the components is .4 g/cm (1400 kg/m2), what particle size should the steroid and excipients be?
Answer #1

Formula: Density = mass/volume

Now, the mass of tablet = 50 mg = 50*10-3 g

i.e. The volume of tablet = 50*10-3 g/(1.4 g/cm3) = 0.0357 cm3

Therefore, the particle size of the tablet = (0.0357 cm3)1/3 = 0.32932 cm

Therefore, for 75*10-6 g steroid, the particle size of the steroid = (75*10-6 g/50*10-3 g)*0.32932 cm = 4.94*10-4 cm

And the particle size of excipient = 0.32932 cm - 4.94*10-4 cm = 0.32882 cm